OpenStudy (anonymous):

A projectile if fired straight upward with an initial velocity of 400 feet per second is modeled by the equation h(t) = - 16t^2 + 400t. How long does it take the projectile to reach the maximum height? 6.25 seconds 2500 seconds 12.5 seconds 25 seconds

OpenStudy (anonymous):

\[h(t) = - 16t^2 + 400t\] max at \[-\frac{b}{2a}=-\frac{400}{2\times (-16)}\]

OpenStudy (anonymous):

I got A..

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\(400\div 32=12.5\)

OpenStudy (anonymous):

Ughhhh....